OpenStudy (anonymous):

Evaluate. 90 • 4 – 14 ÷ 2 + 102 ÷ 5 • 2

OpenStudy (anonymous):

@Brittx.x

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@flower2468

OpenStudy (anonymous):

because the last time the person gave me the answer.

OpenStudy (jameshorton):

110

OpenStudy (jameshorton):

cause 90x4 360

OpenStudy (jameshorton):

360-14

OpenStudy (jameshorton):

346

OpenStudy (flower2468):

90*4=360,so now it is 360-14/2+102/5*2, 14/2=7, and no wit is,360-7+102/5*2, 102/5=20.4, so now it is, 360-7+20.4*2, 20.4*2, 20.4*2=40.8, so now it is, 360-7+40.8, 360-7=299, so now it is, 299+40.8=339.8, I think that is correct

OpenStudy (anonymous):

339.8 is correct?

OpenStudy (anonymous):

ok

OpenStudy (flower2468):

I think ask someone else too

OpenStudy (anonymous):

can u help me with more?

OpenStudy (anonymous):

ok

OpenStudy (flower2468):

I am sorry I have to go

oh):